OpenStudy (anonymous):

a= 1/2 bh solve for h

OpenStudy (anonymous):

a) b=2A/H B)H=2A/B C)H=A/2B D)h=b/2a

OpenStudy (mathstudent55):

h is on the right side. h is being multiplied by 1/2 and by b. You need to multiply both sides by the reciprocals of what h is being multiplied by. That means, multiply both sides by 2 and by 1/b to isolate h.

OpenStudy (anonymous):

& thanks

OpenStudy (mathstudent55):

You're welcome.